If there is 16 side dishes and you have to choose 3, how many different options are there?
Answer by Alan3354(69443)

If there is 16 side dishes and you have to choose 3, how many different options are there?
-------------
The 1st choice is 1 of 16, then 15, then 14
= 16*15*14 = 3360
---
But, choosing A, B & C is the same as B, C & A
There are 3*2*1 = 6 ways to choose A, B & C
3360/6 = 560 different options